OpenStudy (anonymous):

two minor league basebal players got a total of 396 hits. washingtoon has 8 more hits than sanchez. find the number of hits for each player

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

"two minor league basebal players got a total of 396 hits" means that x+y = 396 where x = # of hits washingtoon has y = # of hits sanchez has

jimthompson5910 (jim_thompson5910):

"washingtoon has 8 more hits than sanchez" translates to x = y+8

jimthompson5910 (jim_thompson5910):

x+y = 396 y+8+y = 396 ... plug in x = y+8 2y+8 = 396 ... ... ... y = ???

OpenStudy (anonymous):

so how do i find out how many each has

jimthompson5910 (jim_thompson5910):

by solving 2y+8 = 396 for y

OpenStudy (anonymous):

how do u solve it im in 8th grade

jimthompson5910 (jim_thompson5910):

how do you undo that +8

OpenStudy (anonymous):

you subtract it

jimthompson5910 (jim_thompson5910):

subtract 8 from both sides, good

jimthompson5910 (jim_thompson5910):

so 2y = ???

OpenStudy (anonymous):

396

jimthompson5910 (jim_thompson5910):

396-8 = ??

OpenStudy (anonymous):

388

jimthompson5910 (jim_thompson5910):

so 2y = 388

jimthompson5910 (jim_thompson5910):

then you divide both sides by 2 to get 2y = 388 2y/2 = 388/2 y = 194

OpenStudy (anonymous):

so is 194 for washing or sanchez

jimthompson5910 (jim_thompson5910):

go back up and tell me what y is equal to

OpenStudy (anonymous):

so washington

jimthompson5910 (jim_thompson5910):

go back up where I define x and y

OpenStudy (anonymous):

oops sanchez

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

then plug y = 194 into x = y+8 to find x

OpenStudy (anonymous):

202

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so 202 for washington

jimthompson5910 (jim_thompson5910):

that is correct